OpenStudy (anonymous):

Write the expression in standard form: 3/(3-12i)

OpenStudy (zzr0ck3r):

multiply top and bottom by 3+12i 3(3+12i)/(9-144)

OpenStudy (zzr0ck3r):

does that make sense?

OpenStudy (anonymous):

yeah how do you put it in standard form though?

OpenStudy (zzr0ck3r):

=(9+36i)/(-135) = (-9/135)-(36i/135)

OpenStudy (zzr0ck3r):

simplify the fractions

OpenStudy (anonymous):

right got it so it would be...

OpenStudy (anonymous):

-1/17- 4i/17?

OpenStudy (zzr0ck3r):

-1/15-4i/15?

OpenStudy (zzr0ck3r):

135/9=15?

OpenStudy (zzr0ck3r):

all good:)

OpenStudy (zzr0ck3r):

wait you were right

OpenStudy (anonymous):

huh?

OpenStudy (zzr0ck3r):

3/(3-12i) = (9+36i)/(9+144) = 9/154 + 36/144 = 1/17+4i/17

OpenStudy (zzr0ck3r):

sorry I didn't multiply by a negative the first time

OpenStudy (anonymous):

right all good man thanks!

OpenStudy (zzr0ck3r):

ok good you see what I did?

OpenStudy (anonymous):

Yup